OpenStudy (anonymous):

a dealer bought an article for $7, sold it for $8, bought it back for $9. How much profit did he make?

OpenStudy (anonymous):

hint: when sold + when bought -

OpenStudy (anonymous):

loss of one...

OpenStudy (anonymous):

-8 ?

OpenStudy (anonymous):

At first he makes a profit of $1.......... but late makes a loss of $2..... as a result he makes a loss of $1

OpenStudy (anonymous):

no u have $10-7=3 3+8=11 11-9=2 10-2=8 => loss is 8